OpenStudy (anonymous):

7/10+5/8+7/20= and how

OpenStudy (anonymous):

67/40

OpenStudy (anonymous):

the least common den. is 40 \[\frac{ 7 }{ 10 }*\frac{ 4 }{ 4 }+\frac{ 5 }{ 8 }*\frac{ 5 }{ 5 }+\frac{ 7 }{ 20 }*\frac{ 2 }{ 2 }\] \[\frac{ 28 }{ 40 }+\frac{ 25 }{ 40 }+\frac{ 14 }{ 40 }\] \[\frac{ 28 + 25+14 }{ 40 }=\frac{67}{40}\]

OpenStudy (anonymous):

How do you get a proper fraction from there?

OpenStudy (anonymous):

writing the multiples of 10,8 and 20 10=10 20 30 (40) 50 60 it takes four(4) of the tens to get 40 8=8 16 24 32 (40) 48 it takes five(5) 0f the eights to get 40 20=20 (40) 60 80 it takes two(2) of the twenty to get 40 comparing all the three lowest common multiple(l.c.m) is 40 which is common in all the original equation is; 7/10+5/8+7/20 since the l.c.m =40 we divide through by 40 ((7*4)+(5*5)+(7*2)) /40 this gives (28+25+14)/40 the final answer is 67/40